OpenStudy (anonymous):

find the midpoint of AB whose endpoints are A(37,-70)and B(53,-102)

OpenStudy (anonymous):

the midpoint is like the average of the x's and the y's\[x _{m}=\frac{37+53}{2}=45\]\[y _{m}=\frac{-70-102}{2}=-86\]the midpoint is (45, -86)

OpenStudy (anonymous):

isn't it 86 not -86

OpenStudy (anonymous):

it should be -86 based on the negative y-coordinates of the given endpoints

OpenStudy (anonymous):

here is a rough sketch; it shows the negative y-coordinate of the midpoint